# Multivariable Behavioral Change Modeling of Epidemics in the Presence of Undetected Infections

### v2

@@ Abstract
Epidemic models are invaluable tools to understand and implement strategies to control the spread of infectious diseases, as well as to inform public health policies and resource allocation. However, current modeling approaches have limitations that reduce their practical utility, such as the exclusion of human behavioral change in response to the epidemic or ignoring the presence of undetected infectious individuals in the population. These limitations became particularly evident during the COVID-19 pandemic, underscoring the need for more accurate and informative models. Motivated by these challenges, we develop a novel Bayesian epidemic modeling framework to better capture the complexities of disease spread by incorporating behavioral responses and undetected infections. In particular, our framework makes three contributions: 1) leveraging additional data on hospitalizations and deaths in modeling the disease dynamics, 2) accounting for data uncertainty arising from the large presence of asymptomatic and undetected infections, and 3) allowing the population behavioral change to be dynamically influenced by multiple data sources (cases and deaths). We thoroughly investigate the properties of the proposed model via simulation, and illustrate its utility on COVID-19 data from Montréal and Miami.
@@ 1 Introduction
Epidemic models are essential tools for understanding transmission dynamics and informing public health responses. The foundational SIR compartmental model (Kermack and McKendrick,, 1927), describes disease dynamics by classifying individuals as Susceptible, Infectious, or Removed and using parameters to describe the rates of flow between compartments. These models can be solved deterministically with ordinary differential equations, but stochastic Bayesian implementations are often preferred due to their ability to incorporate prior knowledge and impute missing data through data augmentation (O’Neill and Roberts,, 1999; Lekone and Finkenstädt,, 2006). Most SIR models neglect the role of human behavior changes in shaping transmission dynamics, however, this is a critical limitation that reduces their practical utility and accuracy, a fact that was exposed by the COVID-19 pandemic.
@@
Several behavioral change SIR models have been proposed in the deterministic literature (see Funk et al., (2010) for a comprehensive review), however, these deterministic approaches do not perform parameter estimation using observed data. Recently a stochastic Bayesian behavioral change modeling framework was developed where transmission is dynamically modified by a so-called ‘alarm’ function, which captures the time-varying changes in the population engaging in protective behaviors Ward et al., (2023, 2025). This framework allows for inference on transmission and behavioral parameters and was shown to provide superior model fit to outbreaks of COVID-19, Ebola, and Foot and Mouth Disease, compared to models ignoring behavioral change. Despite this success, these models rely on simplifying assumptions: the use of a standard SIR structure and a behavioral alarm that depends on only one measure of disease severity (incidence or prevalence).
@@
Given the increased availability of hospitalization and death data during an outbreak, there is strong motivation to develop data-driven models that account for these disease states. However, most fully stochastic Bayesian epidemic models are limited to simple SIR or SEIR compartmental structures to reduce computational complexity (Andersson and Britton,, 2012). While expanded compartmental structures (e.g., SIHVR to include hospitalizations and vaccinations (Gibson et al.,, 2023)) have been explored in computationally efficient, partially stochastic models, these approaches neglect the important uncertainty in unobserved compartments by modeling transitions deterministically.
@@
Throughout the COVID-19 pandemic, modeling has been complicated by the large presence of undetected infections (Menachemi,, 2020; Irons and Raftery,, 2021). Existing deterministic or partially stochastic models have addressed this by adding an undetected compartment (Melis and Littera,, 2021; Bhaduri et al.,, 2022) or using a multiplier on the observed case counts (Hao et al.,, 2020; Ma and Rennert,, 2024). In fully stochastic individual-level models, undetected infections can be estimated using reversible jump Markov Chain Monte Carlo (MCMC) techniques (Forrester et al.,, 2007; Jewell and Roberts,, 2012; Ward et al.,, 2021). However, these approaches are computationally infeasible for population-averaged models in large populations as they require imputing disease status for each individual. Thus, there is a need to develop efficient methodology for incorporating undetected infections in fully stochastic population-averaged models.
@@
Motivated by these challenges, we extend the fully stochastic population-averaged behavioral change model to 1) include hospitalizations and deaths in an SIHRD model, 2) incorporate undetected infections, and 3) use a multivariable alarm to quantify the relative importance of two data sources on behavior. In Section 2, we review the univariable alarm SIR model and present our extended SIHRD model with a multivariable alarm and incorporating undetected infections. In Section 3, the proposed methodology is rigorously investigated through simulations. An analysis of two waves of COVID-19 in Miami, Florida and Montréal, Québec is performed in Section 4 to illustrate the utility and limitations of the proposed modeling approach on real data. Concluding remarks are provided in Section 5.
@@ 2.1 Behavioral Change SIR Model
We start by briefly introducing the discrete time behavioral change SIR model of Ward et al., (2023) which serves as the foundation for our work. Assume a closed population of size $N$ and let $S_{t}$, $I_{t}$, and $R_{t}$ denote the number of individuals in the susceptible, infectious, and removed compartments in the continuous time interval $[t,t+1)$, respectively. Transition vectors $I^{*}_{t}$ and $R^{*}_{t}$ represent the number of individuals entering the indicated compartment in this interval. Transitions are temporally described by
@@
Of primary interest is the transmission probability $\pi_{t}^{(SI)}$, traditionally specified as $\pi_{t}^{(SI)}=1-\exp\left(-\beta\frac{I_{t}}{N}\right)$. This probability is derived from assumptions of an independent Poisson contact process and constant infection probability given a contact, with $\beta$ capturing both the contact rate and infection probability as these are not separately identifiable (Brown et al.,, 2016). Behavioral change is incorporated by allowing the transmission rate, $\beta$, to be modified by a time-varying level of alarm, $a_{t}$, in the population as
@@
The alarm is constrained to be in the interval $[0,1]$, and is interpreted as the proportional reduction in transmission due to the alarm in the population. When $a_{t}=0$, the population is in its natural ‘unalarmed’ state, and transmission is described only by $\beta$. When $a_{t}=1$, the population is in its maximal alarmed state and transmission is reduced to zero. In Ward et al., (2023), the alarm was specified as a function of previously observed incidence smoothed over the past $m$ days, i.e., $a_{t}=f\left(I^{*}_{t-1,m}\right)$, where $I^{*}_{t-1,m}=\frac{1}{m}\sum_{i=t-m-1}^{t-1}I^{*}_{i}$ with smoothing parameter $m\in\{1,2,...,t-1\}$. Several functions have been considered to describe the alarm, and we will use the one parameter ‘power’ alarm with growth rate $k>0$
@@
The removal probability is typically specified as $\pi^{(IR)}=1-\exp\left(-\gamma\right)$, which derives from an exponentially distributed length of the infectious period. However, this may be a restrictive assumption and without complete data on both $\boldsymbol{I^{*}}$ and $\boldsymbol{R^{*}}$ would require data-augmented MCMC approaches to implement (Lekone and Finkenstädt,, 2006). To incorporate more flexible modeling of the infectious period and avoid data augmentation, we use the infectious duration-dependent (IDD) specification of Ward et al., (2022), which was shown to be computationally efficient and to improve estimation relative to traditional data augmentation. Let $I_{wt}$ be the number of individuals on day $w$ of the infectious period at time $t$ and $T_{I}$ denote the fixed length of the infectious period. The IDD transmission probability is
@@
where $f(w)$ is a function defining the weighted contribution of infectious individuals depending on how long they’ve been infectious. In our implementation, we use a logistic decay function with a maximum curve value of 1, $f(w)=1/[1+\exp\{\nu(w-w_{0})\}]$, where $w_{0}$ provides the inflection point of the curve, and $\nu>0$ provides the decay rate. As the length of the infectious period is fixed, removal times are fully specified as $R^{*}_{t+T_{I}}=I^{*}_{t}$, for all $t$ and we no longer need to estimate the removal probability $\pi^{(IR)}$. Instead, the parameters of the logistic decay curve are used to quantify the length of time infectious individuals are transmitting the disease. The logistic decay curve assumes peak transmissibility at the start of the infectious period, which may not always be realistic. Alternative IDD curves can allow later peaks, but Ward et al., (2022) found minimal changes in overall transmission estimation between various curves and the logistic decay curve was found to have the highest MCMC efficiency.
@@ 2.2 Multivariable Alarm in an SIHRD model with Undetected Infections
The second extension to the model is to allow for the presence of undetected infections. This captures any individual who did not have a confirmed positive test reported to health officials. Undetected infections are incorporated by splitting the infectious compartment into two parts: observed (detected) cases, $C_{t}$, and undetected infectious individuals, $U_{t}$, such that $I_{t}=C_{t}+U_{t}$. Correspondingly, $I_{t}^{*}=C_{t}^{*}+U_{t}^{*}$, where $C_{t}^{*}$ are the observed cases counts and $U_{t}^{*}$ is unobserved. We assume $C_{t}^{*}\sim Binom\left(I_{t}^{*},\pi^{detect}\right)$, where $\pi^{detect}$ is the probability of an infection being detected. In many cases, there may be data from seroprevalence studies available to inform a strong prior on this detection probability. The proposed SIHRD model with undetected infections is visualized in Web Appendix A.
Lastly, we consider extending the alarm function of Ward et al., (2023) to allow multiple data sources to inform population behavior. The transmission probability $\pi_{t}^{(SI)}$ remains unchanged from Equation 1, but we now consider the alarm to be a function of observed cases and deaths, i.e., $a_{t}=f(C_{t-1,m}^{*},D^{*}_{t-1,m})$. While theoretically the alarm could also depend on other factors, we found that when several highly correlated data sources were used (e.g., cases, hospitalizations, and deaths), a lack of identifiability inhibited estimation. Cases and deaths were chosen as those are most commonly available and because they are furthest apart in the disease process which reduces correlation. In a model with undetected infections the alarm could be specified as a function of the true incidence, $I_{t-1,m}^{*}$, however, as this is generally unknown to the population during an outbreak we find it more realistic to assume that behavior is only informed by observed data. We define the multivariable alarm as
@@ 2.3 Reproductive Number Calculation
Often, the primary inferential interest in compartmental epidemic models is the reproductive number as it quantifies the magnitude of disease spread in a population. We consider the time-varying “effective” reproductive number, $\mathcal{R}_{0}(t)$, defined as the expected number of secondary infections caused by a single infected individual in the current population at time $t$, as our interest is in incorporating behavioral change leading to changing transmission over time. As shown in Ward et al., (2022), this can be generally derived as an expectation from the chain binomial SIR model presented in Section 2.1 as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\pi_{0j}^{(SI)}\left(1-\pi^{(IR)}\right)^{j-t}$, where for a single individual infected at time $t$ $\pi_{0j}^{(SI)}$ denotes the probability of transmission from that individual at time $j$. Intuitively, the formulation is derived from assuming that on any given day, the expected number of secondary infections is $S_{j}\times\pi_{0j}^{(SI)}$, and then over the course of the infection this quantity is weighted by the probability that the individual is still infectious (i.e., has not become removed). For the IDD transmissibility specification we are using to implement the SIR model, the duration of the infection is fixed so this simplifies to $\mathcal{R}_{0}(t)=\sum_{j=t}^{t+T_{I}}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{f(j-t)}{N}\Big\}\Big]$. This formulation can be directly extended to the SIHRD compartmental structure, however, we must be sure to account for the multinomial transition out of the Infectious compartment as according to the model individuals can stop transmitting the disease when they recover or when they become hospitalized. Thus, the effective reproductive number is calculated as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{1}{N}\Big\}\Big]\left(1-\pi^{(IH)}-\pi^{(IR)}\right)^{j-t}$.

### v3

@@ Abstract
Epidemic models are invaluable tools to understand and implement strategies to control the spread of infectious diseases, as well as to inform public health policies and resource allocation. However, current modeling approaches have limitations that reduce their practical utility, such as the exclusion of human behavioral change in response to the epidemic or ignoring the presence of undetected infectious individuals in the population. These limitations became particularly evident during the COVID-19 pandemic, underscoring the need for more accurate and informative models. To address these challenges, we develop a novel Bayesian epidemic modeling framework to better capture the complexities of disease spread by incorporating behavioral responses and undetected infections. In particular, our framework makes three contributions: 1) leveraging additional data on hospitalizations and deaths in modeling the disease dynamics, 2) accounting for data uncertainty arising from the large presence of asymptomatic and undetected infections, and 3) allowing the population behavioral change to be dynamically influenced by multiple data sources (cases and deaths). We thoroughly investigate the properties of the proposed model via simulation, and illustrate its utility on COVID-19 data from Montréal and Miami.
@@ 1 Introduction
Epidemic models are essential tools for understanding transmission dynamics and informing public health responses. Infectious disease transmission is influenced by many factors, including the infectiousness of the disease, the susceptibility of the population, and the behavior of the population in response to the outbreak. Behavioral change could include city- or state-wide lockdown policies, school closures, or personal changes such as voluntary masking and can result in a delayed or reduced peak or multiple waves. Despite the fact that human behavior drives disease transmissionarthur2017contact, many traditional modeling approaches do not account for the impact of human behavior changing transmission dynamics, a fact that was exposed by the COVID-19 pandemic. The incorporation of behavioral change in epidemic models can help public health officials better understand disease spread and the impact of interventions or messaging, ultimately leading to better outbreak control(osi2025simultaneous).
Compartmental epidemic models describe infectious disease dynamics by dividing the population into epidemiologically important disease states. The foundational compartmental model of Kermack and McKendrick(kermack1927) is the Susceptible, Infectious, or Removed (SIR) model, which specifies that individuals start in the susceptible compartment, become infected through contact with infectious individuals, and become removed when they recover with immunity or die from the disease. Parameters describe the flow between compartments, with the transition from susceptible to infectious being of primary interest as this captures disease transmission. These models can be specified deterministically through a series of ordinary differential equations, or incorporate stochasticity by probabilistically describing the parameters or the transitions between compartments. Stochastic models are more realistic representations of the spread of an infectious pathogen, and incorporating this uncertainty is particularly useful when describing disease spread at the start of an epidemic or in small populations roberts2015nine. Both the deterministic and stochastic approaches can be used to simulate epidemic trajectories for the purposes of comparing potential behavioral dynamics or interventions, or can be used to make inference on parameters using observed data in either a frequentist or Bayesian context. We consider the stochastic Bayesian implementation where the primary goal is making inference on transmission and behavioral parameters using motivated by a real observed epidemic. The Bayesian framework is advantageous in this setting due to its ability to provide uncertainty quantification that naturally accounts for parameter estimation, incorporate prior knowledge on some parameters, and impute missing data through data augmentation (o1999bayesian; lekone).
@@
Several behavioral change SIR models have been proposed prior to the occurrence of the COVID-19 pandemic (see Funk et al funk2010modelling and Verelst et al verelst2016behavioural for reviews of behavioral change modeling approaches published through 2015). The onset of the pandemic brought increased attention towards epidemic modeling, however, a recent review found only 38% of models attempted to incorporate human behavior lee2025review. Additionally a large majority of the behavioral change literature both pre- and post-COVID-19 focused on model-specification and simulation, and not on performing parameter estimation using observed data funk2010modelling; verelst2016behavioural; lee2025review. In many cases, this is due to highly specified behavioral dynamics that are not supported by currently available data and therefore the associated behavioral parameters are difficult to estimate. Recently a stochastic Bayesian behavioral change modeling framework was developed where transmission is dynamically modified by a so-called ‘alarm’ function, which captures time-varying changes in transmission that may be due to the population engaging in protective behaviors ward2023bayesian; ward2025framework. This framework allows for inference on transmission and behavioral parameters using observed data. It has also been shown to provide superior model fit to outbreaks of COVID-19, Ebola, and Foot and Mouth Disease, compared to models ignoring behavioral change. Despite this success, these models rely on simplifying assumptions: the use of a standard SIR structure and a behavioral alarm that depends on only one measure of disease severity (incidence or prevalence).
@@
Beyond behavioral change models, Bayesian compartmental epidemic models are often simplified to reduce computational complexity (andersson2012stochastic). Most fully stochastic Bayesian compartmental models include only SIR compartments (ward2023bayesian; gu2022bayesian) or are at most expanded to include a latent exposed compartment for an SEIR structure (deardon2010; brownReproductive; ward2021individual). Expanded compartmental structures have been considered in partially stochastic modeling approaches, where transmission and other parameters are estimated using Bayesian inference, but all compartment transitions themselves are deterministic and specified using differential equations. As examples of the partially stochastic approach in application to COVID-19, Self et al self2022bayesian include hospitalizations and vaccinations in an SIHVR model and Gibson et al gibson2023real include a compartment for deaths in an SEIRD model. These hybrid approaches are computationally efficient as unobserved compartments, (e.g., Recovered), do not need to be imputed. However, by definition they do not incorporate important uncertainty in the unobserved compartment membership over time. Given the wide availability of hospitalization and death data on COVID-19, there is strong motivation to develop a fully stochastic modeling framework to explicitly account for these additional disease states and provide a more accurate representation of the disease process.
@@
Throughout the COVID-19 pandemic, modeling has been complicated by the large presence of undetected infections, either due to lack of testing availability, asymptomatic infection, or false negative test results (menachemi2020population; irons2021estimating). Existing models have addressed this by adding an undetected compartment (melis2021undetected; bhaduri2022extending), using a multiplier on the observed case counts (hao2020reconstruction; ma2024epidemiological), or using a capture-recapture framework to estimate a lower bound on the number of undetected infections (bohning2020estimating). In individual-level models, undetected infections can be estimated using reversible jump Markov Chain Monte Carlo (MCMC) techniques (forrester2007bayesian; jewell2012enhancing; ward2021individual). However, this approach is computationally infeasible for population-averaged models in large populations as it would require imputing the infection status for each individual.
@@
Motivated by these challenges, we extend the fully stochastic population-averaged behavioral change model to 1) include hospitalizations and deaths in an SIHRD model, 2) incorporate undetected infections, and 3) use a multivariable alarm to quantify the relative importance of two data sources on behavior. In combination, this yields a flexible behavioral change modeling framework that can provide important and accurate inference on both transmission and behavioral dynamics. In Section 2, we review the univariable alarm SIR model and present our behavioral change SIHRD framework with a multivariable alarm and incorporating undetected infections. In Section 3, the proposed methodology is rigorously investigated through simulations. An analysis of two waves of COVID-19 in Miami, Florida and Montréal, Québec is performed in Section 4 to illustrate the utility and limitations of the proposed modeling approach on observed data. Concluding remarks are provided in Section 5.
@@ 2.1 Behavioral Change SIR Model
We start by briefly introducing the discrete time behavioral change SIR model of Ward et alward2023bayesian that serves as the foundation for our work. Assume a closed population of size $N$ and let $S_{t}$, $I_{t}$, and $R_{t}$ denote the number of individuals in the susceptible, infectious, and removed compartments in the continuous time interval $[t,t+1)$, respectively. Transition vectors $I^{*}_{t}$ and $R^{*}_{t}$ represent the number of individuals entering the indicated compartment in this interval. Transitions are temporally described by
@@
Of primary interest is the transmission probability $\pi_{t}^{(SI)}$, traditionally specified as $\pi_{t}^{(SI)}=1-\exp\left(-\beta\frac{I_{t}}{N}\right)$. This probability is derived from assumptions of an independent Poisson contact process and constant infection probability given a contact, where $\beta$ captures both the contact rate and infection probability as these are not separately identifiable (brownReproductive). Behavioral change is incorporated by allowing the transmission rate, $\beta$, to be modified by a time-varying level of alarm, $a_{t}$, in the population as
@@
The alarm is constrained to be in the interval $[0,1]$, and is interpreted as the proportional reduction in transmission due to the alarm in the population. When $a_{t}=0$, the population is in its natural ‘unalarmed’ state, and transmission is described only by $\beta$. When $a_{t}=1$, the population is in its maximal alarmed state and transmission is reduced to zero. In Ward et alward2023bayesian, the alarm was specified as a function of previously observed incidence smoothed over the past $m$ days, i.e., $a_{t}=f\left(I^{*}_{t-1,m}\right)$, where $I^{*}_{t-1,m}=\frac{1}{m}\sum_{i=t-m-1}^{t-1}I^{*}_{i}$ with smoothing parameter $m\in\{1,2,...,t-1\}$. Several functions have been considered to describe the alarm, and we will use the one parameter ‘power’ alarm with growth rate $k>0$
@@
The removal probability is typically specified as $\pi^{(IR)}=1-\exp\left(-\gamma\right)$, which derives from an exponentially distributed length of the infectious period. However, this may be a restrictive assumption and without complete data on both $\boldsymbol{I^{*}}$ and $\boldsymbol{R^{*}}$ would require data-augmented MCMC approaches to implement (lekone). To incorporate more flexible modeling of the infectious period and avoid data augmentation, we use the infectious duration-dependent (IDD) specification of Ward et alward2022idd, which was shown to be computationally efficient and to improve estimation relative to traditional data augmentation. Let $I_{wt}$ be the number of individuals on day $w$ of the infectious period at time $t$ and $T_{I}$ denote the fixed length of the infectious period. The IDD transmission probability is
@@
where $f(w)$ is a function defining the weighted contribution of infectious individuals depending on how long they’ve been infectious. In our implementation, we use a logistic decay function with a maximum curve value of 1, $f(w)=1/[1+\exp\{\nu(w-w_{0})\}]$, where $w_{0}$ provides the inflection point of the curve, and $\nu>0$ provides the decay rate. As the length of the infectious period is fixed, removal times are fully specified as $R^{*}_{t+T_{I}}=I^{*}_{t}$, for all $t$. Therefore, we no longer need to estimate the removal probability $\pi^{(IR)}$. Instead, the parameters of the logistic decay curve are used to quantify the length of time infectious individuals are transmitting the disease. The logistic decay curve assumes peak transmissibility at the start of the infectious period, which may not always be realistic. Alternative IDD curves can allow later peaks, but Ward et alward2022idd found minimal changes in overall transmission estimation between various curves and the logistic decay curve was found to have the highest MCMC efficiency, defined as the number of effective samples of the reproductive number per unit time.
@@ 2.2 Multivariable Alarm in an SIHRD model with Undetected Infections
The second extension to the model is to allow for the presence of undetected infections. This captures any individual who did not have a confirmed positive test reported to health officials. Undetected infections are incorporated by splitting the infectious compartment into two parts: observed (detected) cases, $C_{t}$, and undetected infectious individuals, $U_{t}$, such that $I_{t}=C_{t}+U_{t}$. Correspondingly, $I_{t}^{*}=C_{t}^{*}+U_{t}^{*}$, where $C_{t}^{*}$ are the observed cases counts and $U_{t}^{*}$ is unobserved. We assume $C_{t}^{*}\sim Binom\left(I_{t}^{*},\pi^{detect}\right)$, where $\pi^{detect}$ is the probability of an infection being detected, i.e., the case ascertainment rate. In many cases, there may be data from seroprevalence studies available to inform a strong prior on this detection probability. The proposed SIHRD model with undetected infections is visualized in Supplemental Figure 1. It’s worth noting that we assume a constant probability of detection over time, however, in reality this may be time-varying due to changes in testing capacity (omori2020changes) or test-seeking behavior (colman2023ascertainment). The assumption of time-invariant case ascertainment was used to better facilitate model identifiability and may be reasonable for shorter time scales such as the six to ten weeks used in our data analyses. We provide further consideration of this assumption and alternative approaches in Section 5.
Lastly, we consider extending the alarm function of Ward et alward2023bayesian to allow multiple data sources to inform population behavior. The transmission probability $\pi_{t}^{(SI)}$ remains unchanged from Equation 1, but we now consider the alarm to be a function of observed cases and deaths, i.e., $a_{t}=f(C_{t-1,m}^{*},D^{*}_{t-1,m})$. While theoretically the alarm could also depend on other factors, we found that when several highly correlated data sources were used (e.g., cases, hospitalizations, and deaths), a lack of identifiability inhibited estimation. Cases and deaths were chosen as those are most commonly available and because they are furthest apart in the disease process which reduces correlation. In a model with undetected infections the alarm could be specified as a function of the true incidence, $I_{t-1,m}^{*}$, however, as this is generally unknown to the population during an outbreak we find it more realistic to assume that behavior is only informed by observed data. We define the multivariable alarm as
@@ 2.3 Reproductive Number Calculation
Often, the primary inferential interest in compartmental epidemic models is the reproductive number as it quantifies the magnitude of disease spread in a population. We consider the time-varying “effective” reproductive number, $\mathcal{R}_{0}(t)$, defined as the expected number of secondary infections caused by a single infected individual in the current population at time $t$, as our interest is in incorporating behavioral change leading to changing transmission over time. As shown in Ward et alward2022idd, this can be generally derived as an expectation from the chain binomial SIR model presented in Section 2.1 as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\pi_{0j}^{(SI)}\left(1-\pi^{(IR)}\right)^{j-t}$, where for a single individual infected at time $t$ $\pi_{0j}^{(SI)}$ denotes the probability of transmission from that individual at time $j$. Intuitively, the formulation is derived from assuming that on any given day, the expected number of secondary infections is $S_{j}\times\pi_{0j}^{(SI)}$, and then over the course of the infection this quantity is weighted by the probability that the individual is still infectious (i.e., has not become removed). For the IDD transmissibility specification we are using to implement the SIR model, the duration of the infection is fixed so this simplifies to $\mathcal{R}_{0}(t)=\sum_{j=t}^{t+T_{I}}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{f(j-t)}{N}\Big\}\Big]$. This formulation can be directly extended to the SIHRD compartmental structure, however, we must be sure to account for the multinomial transition out of the Infectious compartment as according to the model individuals can stop transmitting the disease when they recover or when they become hospitalized. Thus, the effective reproductive number is calculated as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{1}{N}\Big\}\Big]\left(1-\pi^{(IH)}-\pi^{(IR)}\right)^{j-t}$.